OpenStudy (anonymous):

Solve the following equation, answer as a reduced, mixed number. Then place the correct number in the box provided. 4(7x - 1) = 3(2x + 1) - 4(x + 6)

OpenStudy (anonymous):

Okay x=\[-\frac{ 17 }{ 26 }\]

OpenStudy (anonymous):

@Nahstya, you should present how you arrived at the answer. It is against the Code of Conduct to just give answers.

OpenStudy (anonymous):

$$4(7x - 1) = 3(2x + 1) - 4(x + 6)$$ Distribute. $$28x - 4 = 6x + 3 - 4x - 24$$ Simplify. $$28x - 4 = 2x - 21$$ Add 4 to both sides. $$28x = 2x - 17$$ Subtract 2x from both sides. $$26x = -17$$ Divide 26 by both sides. $$x = -\frac{17}{26}$$

OpenStudy (anonymous):

Lol ok? I didn't know....